ELECTRONIC STRUCTURE AND SPECTROSCOPY OF NH^{+}

\underline{G. J. VÁZQUEZ}, and J. M. AMERO, Centro de Ciencias Físicas, Universidad Nacional Autónoma de México (UNAM), 62251 Cuernavaca, México (vaztor@fis.unam.mx).

Ab–initio SCF MRSD–CI electronic structure calculations were carried out on the NH^{+} cation. A basis set of DZ+POL quality augmented with Rydberg and bond functions was employed together with an extensive treatment of electron correlation. More than fifty electronic states of NH^{+} are reported, including doublets, quartets and sextets^{a}. Leading configurations, vertical ionization energies of NH, vertical excitation energies of NH^{+} and potential energy curves are reported. Spectroscopic properties calculated for the five known bound electronic states of NH^{+} (X^{2}Π, a^{4}Σ^{−}, A^{2}Σ^{−}, B^{2}Δ, C^{2}Σ^{+}) are found in good agreement with experiment. We find four new bound electronic states of NH^{+}, namely, three doublets (^{2}Δ, ^{2}Σ^{+}, ^{2}Σ^{+}) and one sextet (^{6}Π). A fifth bound electronic state, a quartet (^{4}Π) which was reported schematically in an early SCF study^{b} is calculated for the first time over a wide range of N–H distances. Adiabatic excitation and ionization energies and spectroscopic constants are also reported for these five states^{c}, as well as vertical de–excitation energies for emission from the novel states to lower lying bound states. The potential energy curve of NH^{2+}(X^{1}Σ^{+}) was also calculated to aid in the discussion of a Rydberg bound state of NH^{+}. It is hoped that the information reported in the present work will be helpful in guiding the laboratory and/or astrophysical search for these species.

aJ. M. Amero, G. J. Vázquez, Int. J. Quantum Chem., 101, 396 (2005).

bH. P. D. Liu, G. J. Verhaegen, J. Chem. Phys., 53, 735 (1970).

cJ. M. Amero, G. J. Vázquez, Int. J. Quantum Chem., 99, 353 (2004).